# for Large-scale Vision-Language Models

### v1

@@ Abstract
Large-scale Vision-Language Models (LVLMs) have significantly advanced with text-aligned vision inputs. They have made remarkable progress in computer vision tasks by aligning text modality with vision inputs. There are also endeavors to incorporate multi-vision sensors beyond RGB, including thermal, depth, and medical X-ray images. However, we observe that current LVLMs view images taken from multi-vision sensors as if they were in the same RGB domain without considering the physical characteristics of multi-vision sensors. They fail to convey the fundamental multi-vision sensor information from the dataset and the corresponding contextual knowledge properly. Consequently, alignment between the information from the actual physical environment and the text is not achieved correctly, making it difficult to answer complex sensor-related questions that consider the physical environment. In this paper, we aim to establish a multi-vision Sensor Perception And Reasoning benchmarK called SPARK that can reduce the fundamental multi-vision sensor information gap between images and multi-vision sensors. We generated 6,248 vision-language test samples automatically to investigate multi-vision sensory perception and multi-vision sensory reasoning on physical sensor knowledge proficiency across different formats, covering different types of sensor-related questions. We utilized these samples to assess ten leading LVLMs. The results showed that most models displayed deficiencies in multi-vision sensory reasoning to varying extents. Codes and data are available at https://github.com/top-yun/SPARK
@@ Introduction
In this paper, to handle the aforementioned challenge, we design the SPARK benchmark to evaluate multi-vision input LVLMs on two fronts: multi-vision perception and multi-vision reasoning. Multi-vision perception pertains to the information needed, which measures the LVLM’s effectiveness in satisfying visual perception needs. Multi-vision reasoning measures the LVLM’s ability to base its responses on fundamental information from the provided sensor knowledge. To be specific, we generated 6,248 vision-language test samples automatically to investigate multi-vision sensory perception and reasoning related to physical sensor knowledge proficiency, covering 6 types of multi-vision sensory instruction tasks across 2 different question-and-answer formats. We used these samples to assess 10 leading large-scale vision language models. The experiment results validate that most LVLMs displayed deficiencies in sensory reasoning to varying extents.
@@
Using accuracy as the evaluation metric for both multi-choice questions and Yes-or-No questions ensures consistency in how we assess the model’s performance. Accuracy, defined as the proportion of correctly answered questions, provides a clear and intuitive measure of how well the model understands the given inputs. The adoption of the multi-choice question based evaluation design facilitates the identification of LVLM’s performance on specific sensor knowledge. By analyzing patterns in the model’s responses, we can gain insights into which types of questions or subject areas the model excels in and which areas require further improvement. This diagnostic capability is essential for guiding future development and refinement of multi-vision sensor models. Additionally, the multi-choice question format supports the development of a more comprehensive evaluation framework by incorporating a diverse range of question types. This diversity ensures that the model is tested across a wide spectrum of scenarios, providing a more holistic assessment of its capabilities. The incorporation of both simple Yes-or-No questions and more complex multi-choice questions ensures that the evaluation covers both basic and advanced aspects of LVLM’s understanding.
@@
Multi-vision perception is the foundational process by which Large Vision-Language Models (LVLMs) analyze images captured by various multi-vision sensors, including RGB, thermal, depth, and X-ray images. This process involves recognizing and interpreting the fundamental elements within each image.
@@ Conclusion
In this study, we focus on evaluating the ability of Large Vision-Language Models (LVLMs) to understand and process multi-vision sensory inputs. As LVLMs are increasingly deployed in real-world applications, their ability to accurately interpret and reason about data from diverse vision sensors has become crucial. To address this, we propose an evaluation benchmark called SPARK, which automatically generates instruction tuning samples aimed at specific physical sensor understanding in various question-and-answer formats. Through extensive experiments, we assess the performance of understanding sensory knowledge in the latest state-of-the-art LVLMs handling multi-vision input. We believe this approach, integrating a sensory knowledge annotated evaluation benchmark paves the way for promising future applications of LVLMs.

### v2

@@ Abstract
Large-scale Vision-Language Models (LVLMs) have significantly advanced with text-aligned vision inputs. They have made remarkable progress in computer vision tasks by aligning text modality with vision inputs. There are also endeavors to incorporate multi-vision sensors beyond RGB, including thermal, depth, and medical X-ray images. However, we observe that current LVLMs view images taken from multi-vision sensors as if they were in the same RGB domain without considering the physical characteristics of multi-vision sensors. They fail to convey the fundamental multi-vision sensor information from the dataset and the corresponding contextual knowledge properly. Consequently, alignment between the information from the actual physical environment and the text is not achieved correctly, making it difficult to answer complex sensor-related questions that consider the physical environment. In this paper, we aim to establish a multi-vision Sensor Perception And Reasoning benchmarK called SPARK that can reduce the fundamental multi-vision sensor information gap between images and multi-vision sensors. We generated 6,248 vision-language test samples to investigate multi-vision sensory perception and multi-vision sensory reasoning on physical sensor knowledge proficiency across different formats, covering different types of sensor-related questions. We utilized these samples to assess ten leading LVLMs. The results showed that most models displayed deficiencies in multi-vision sensory reasoning to varying extents. Codes and data are available at https://github.com/top-yun/SPARK
@@ Introduction
In this paper, to handle the aforementioned challenge, we design the SPARK benchmark to evaluate multi-vision input LVLMs on two fronts: multi-vision perception and multi-vision reasoning. Multi-vision perception pertains to the information needed, which measures the LVLM’s effectiveness in satisfying visual perception needs. Multi-vision reasoning measures the LVLM’s ability to base its responses on fundamental information from the provided sensor knowledge. To be specific, we generated 6,248 vision-language test samples to investigate multi-vision sensory perception and reasoning related to physical sensor knowledge proficiency, covering 6 types of multi-vision sensory instruction tasks across 2 different question-and-answer formats. We used these samples to assess 10 leading large-scale vision language models. The experiment results validate that most LVLMs displayed deficiencies in sensory reasoning to varying extents.
@@
Using accuracy as the evaluation metric for both multi-choice questions and Yes-or-No questions ensures consistency in how we assess the model’s performance. Accuracy, defined as the proportion of correctly answered questions, provides a clear and intuitive measure of how well the model understands the given inputs. The adoption of the multi-choice question based evaluation design supports the development of a more comprehensive evaluation framework. The incorporation of both simple Yes-or-No questions and more complex multi-choice questions ensures that the evaluation covers both basic and advanced aspects of LVLM’s understanding.
@@
Multi-vision perception is the foundational process by which Large Vision-Language Models (LVLMs) analyze images captured by various multi-vision sensors, including RGB, thermal, depth, and X-ray images. This process involves recognizing and interpreting the fundamental elements within each visual input based on cognitive science (Kahneman, Treisman, and Gibbs 1992; Broadbent 2013).
@@
Next, we integrate both visual and textual inputs into GPT-4, guided by meticulously crafted prompts. These prompts are specifically designed to align with each evaluation task, ensuring that the generated questions are both relevant and focused. To further enhance the quality of the benchmark, we introduce an additional filtering step. In the final stages of development, human annotators play a crucial role, selecting the correct answers and categorizing the questions according to their respective evaluation tasks.
@@ Conclusion
In this study, we focus on evaluating the ability of Large Vision-Language Models (LVLMs) to understand and process multi-vision sensory inputs. As LVLMs are increasingly deployed in real-world applications, their ability to accurately interpret and reason about data from diverse vision sensors has become crucial. To address this, we propose an evaluation benchmark called SPARK, which generates instruction tuning samples aimed at specific physical sensor understanding in various question-and-answer formats. Through extensive experiments, we assess the performance of understanding sensory knowledge in the latest state-of-the-art LVLMs handling multi-vision input. We believe this approach, integrating a sensory knowledge annotated evaluation benchmark paves the way for promising future applications of LVLMs.